An impoundment 20 x 12 feet is contaminated by tetrachloroethylene. The contaminated liquid comes in contact with the ground water and a concentration gradient of 1.3 x 10-6 (mol/cm3/cm) is established. Use Fick's law to determine the diffusive flux. Assuming diffusion is the only driving force, how many pounds of tetrachloroethylene will move toward the ground water in one hour?
Answer #1

Solution A20x12 ff? = 240x (0-3048 m)* A = 22297 m2 Dag = 8.2x 10-6 cał/sec VA= 8.2x7-64 82X10 mx 1.3x10-6 mol sec JA = 10.66